OpenStudy (anonymous):

What is the volume of a sphere if it has a radius of 3 centimeters?

OpenStudy (kropot72):

Do you know the formula for the volume of a sphere?

OpenStudy (anonymous):

nope? sorry :(

OpenStudy (reemii):

these formulas are easy to find with google. i just found it : \(V = \frac43\pi r^3\).

OpenStudy (kropot72):

\[volume\ of\ a\ sphere=\frac{4}{3} \pi r ^{3}\] Just plug the value of r into the formula.